Let

ƒ(x) = 𝑥1/(1−x) .

Make tables of values of ƒ at values of x that approach x = 1 from above and below. Does ƒ(x)

appear to have a limit as x approaches 1? If so, what is it? If not, why not?


1
Expert's answer
2021-10-04T16:21:58-0400

Let "y=f(x)=\dfrac{x+1}{1-x}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & y \\ \hline
 0.9 & 19 \\
0.99 & 199 \\
0.999 & 1999 \\
0.9999 & 19999 \\
1.0001 & -20001 \\
1.001 & -2001 \\
1.01 & -201 \\
1.1 & -21 \\
\end{array}"


"\lim\limits_{x\to1^-}f(x)=\infin"

"\lim\limits_{x\to1^+}f(x)=-\infin"

Since "\lim\limits_{x\to1^-}f(x)\not=\lim\limits_{x\to1^+}f(x)," then "\lim\limits_{x\to1}f(x)" does not exist.